An archer shoots an arrow from a height 4.2 m vertical height from the ground with a speed 40m/s and at an angle 30° to horizontal level Determine the horizontal distance R' covered by the arrow when it hits the ground.(Let g=10m/s²)
Ans 84√3
1
Expert's answer
2020-01-24T05:18:00-0500
"h+v\sin{30}t-0.5gt^2=0"

"4.2+40\sin{30}t-0.5(10)t^2=0"

"t=4.2\ s"

So,


"R=v\cos{30}t=40\cos{30}(4.2)=84\sqrt{3}\ m"